# EPOCHS IV: SED Modelling Assumptions and their impact on the Stellar Mass Function at using PEARLS and public JWST observations

### v1

@@ 5.4 Fitting the GSMF
The Schechter function, which consists of a power-law with an exponential fall-off, is parameterised by 3 parameters: , M, and . controls the slope of the low-mass end of the SMF, while M gives the turnover point at which the function turns from a power-law to an exponential fall-off. controls the overall scaling of the SMF, and is the only parameter that has been confidently shown to evolve with cosmic time (Popesso et al., 2023). Stellar mass functions are sometimes split into star-forming and quiescent, with a division in specific star formation rate (sSFR). SMFs for quiescent galaxies are often parameterised by double Schechter functions (Weigel et al., 2016).
Before JWST the mass function was constrained up to using HST WFC3/IR, Spitzer/IRAC and ground-based data (Visible and Infrared Survey Telescope for Astronomy (VISTA), United Kingdom Infrared Telescope (UKIRT)) observations of the HUDF, CANDELS, Cluster Lensing and Supernova survey with Hubble (CLASH) and Hubble Frontier Fields (HFF) surveys (Stark et al., 2013; Bradač et al., 2014; Duncan et al., 2014; Bouwens et al., 2016; Song et al., 2016; Stefanon et al., 2017; Bhatawdekar et al., 2019; Kikuchihara et al., 2020). These studies generally agree that the mass function steepens out to , with the low-mass slope, , . Duncan et al. (2014); Bhatawdekar et al. (2019); Stefanon et al. (2021) find that the stellar mass to halo mass ratio in galaxies (M-M) shows no evolution over despite a 3 dex increase in overall stellar mass, suggesting that stellar and halo mass grew together during reionization. Pre-JWST, the accuracy of stellar mass functions at these redshifts is limited by the uncertainty in stellar mass estimates from Spitzer data, where the broadband photometry can be affected by strong nebular emission causing scatter in stellar mass estimates, which is compounded by the limited number of spectroscopically confirmed galaxies at these redshifts (Roberts-Borsani et al., 2016; Strait et al., 2020). Previously, measurements of the mass function above have relied on UV-selected samples of galaxies, with mass indirectly inferred through a calibrated L-M relationship (Song et al., 2016; Harikane et al., 2016) rather than direct measurements of the rest-frame optical wavelengths. Photometric mass-selected samples above 2 all have significant mass uncertainties at higher redshift (Retzlaff et al., 2010; Laigle et al., 2016; Weaver et al., 2022),. Many studies combine ground and space-based observations in order to probe the entire mass function, although combining these datasets robustly is difficult due to systematics between different surveys, including differing selection functions, SED fitting techniques, detection bands, and survey depths.
@@
For our highest redshift bin (11.5 13.5), we do not attempt to fit the mass function for a number of reasons; firstly given the low number of available data point points a meaningful fit is not possible without fixing multiple parameters, and secondly due to the increasing uncertainties in stellar mass estimates and reliability of candidates at these redshifts. Furthermore the stellar mass estimates are less reliable when there is less rest-frame optical photometry available during SED fitting. Our JAGUAR simulations also suggest that this bin has high contamination at lower mass, with many galaxies below 10 M having 50% average contamination across the fields they are detectable in.
We use wide uniform priors on , M and , with and in our MCMC fitting procedure and run until it has converged before taking independent draws from the posterior. We calculate the median posterior, the uncertainty from the 16th/84th percentiles as well as the maximum likelihood draw.
@@
Figure 9 shows the covariance of the Schechter parameters for each GSMF fit. A range of parameters are able to fit our GSMF estimates, demonstrating the highly covariant nature of the Schechter parameters. We note that the confidence intervals for the turnover mass, M, and the normalization, , appear somewhat unconstrained at the high and low ends respectively. The lack of constraint on M is a consequence of the wide stellar mass bins used and consequential low number of overall mass bins. Wherever we find M11.5, the GSMF is fit by a pure power-law only, and there is no reason to continually extend the prior outside the range where we have available data to constrain M.
@@
| Redshift Bin | Med() | Mpc) | Comp (%) | Cont (%) | N | |
| 6.5 7.5 | 8.07 | 31 | 23 | 273 | ||
@@ 5.6 Stellar Mass Density
We perform the integration given in Equation 5, where is the best-fitting Schechter function fit for each redshift bin. We make use of the quad function from SciPy (Virtanen et al., 2020), integrating between 10 M and 10 M. We compare the integration of both the maximum likelihood fit as well as the median posterior fit. For comparison we also integrate the best-fitting Schechter functions from the literature in each redshift bin, converting to a Kroupa (2001) IMF if necessary.
@@ 6.1 Massive Galaxies in the Early Universe
Numerous studies have reported an excess of galaxies with high inferred stellar mass at high-redshift (Labbé et al., 2023; Endsley et al., 2023b; Akins et al., 2023; Xiao et al., 2023). If galaxies of the inferred masses do exist in the number densities implied by these studies, this could represent a challenge to CDM cosmology or our understanding of high-redshift galaxy astrophysics, given the available timescale and available gas reservoirs (Labbé et al., 2023; Lovell et al., 2023; Boylan-Kolchin, 2023).
@@
In order to test whether our derived stellar mass and redshift estimates are in tension with CDM we use the Extreme-Value Statistics (EVS) approach, presented in Lovell et al. (2023) and available online666github.com/christopherlovell/evstats. We follow the methodology of Lovell et al. (2023) throughout this section. Given the total available area of our survey (187 arcmin) we have computed the halo distribution PDF and corresponding stellar mass PDF.
A universal baryon fraction of 0.16 is assumed based on cosmological results (Planck Collaboration et al., 2016), and we compute the EVS limits for a range of stellar fractions (fraction of baryons which form stars). In Figure 12 we show the stellar fraction upper limit of unity with a dashed line, as well as the more realistic predictions for stellar fraction, following a simple truncated lognormal PDF with , with a dotted line and shaded contours. Overlaid on Figure 12 we show the SED-fitting derived redshifts and corrected stellar masses for our sample. We correct our derived stellar masses for Eddington bias (Eddington, 1913) following . is the stellar mass estimate, is the uncertainty in the stellar mass taken from the posterior PDF, and is the local slope of the halo mass function. Jespersen et al. (2024) looks at the impact of cosmic variance on the predictions of EVS, as this is not accounted for in the method of Lovell et al. (2023) we use here. Given that our sample consists of multiple widely-separated fields, we do not believe cosmic variance will play a significant role on this EVS analysis.
@@
No galaxies in our sample fall above the limit imposed by a stellar fraction of unity, shown with a black line, which would require more stellar mass than the available baryons to form stars. The majority of the EPOCHS v1 sample are not in tension with the EVS limits and are shown with greyscale hexbins on Figure 12, shaded according to number density within the bin. For galaxies with redshift and stellar mass estimates which place them in possible tension with the EVS limits assuming a realistic stellar fraction (more than 1.5 the mean value of the lognormal distribution, shown by the dashed-dot line in Figure 12) we overlay the individual mass measurements as points on Figure 12. For each of these galaxies we show the ‘fiducial’ Bagpipes photo- and stellar mass, as well as the maximum and minimum stellar mass estimate for that galaxy in our Bagpipes and Prospector fitting, including the modified IMF results we discuss in § 4.5. We link individual mass estimates for the same object with dotted lines. In the majority of cases, whilst the maximum stellar mass estimate may suggest a possible tension, the minimum and often the fiducial mass estimates are not in tension.
@@
It is worth noting that all galaxies which show a possible tension at redshift in this figure are classed as ‘little red dots’ (LRDs), which are discussed further in § 5.2.3 and potentially contain a significant AGN component we do not account for. It is clear from Figure 12 that none of our candidates are in tension with the predictions of CDM, as seen by the lack of sources above the solid black line, however a number of objects do require very high stellar fractions at these redshifts. The galaxies which require the highest star formation efficiencies seem to be found at , rather than the higher redshift probed in this study. This has also been found for HST-dark galaxies observed as part of FRESCO, where the galaxies with highest implied stellar mass densities are between (Xiao et al., 2023).
@@
Wang et al. (2024) have shown that stellar masses of high mass galaxies are typically overestimated by 0.4 dex when m rest-frame emission is not used in SED fitting, which requires MIRI observations in this redshift regime. As MIRI observations are not used in this study, the degeneracies in age-attenuation and the relative contributions of strong emission lines and the stellar continuum observed by Wang et al. (2024) are not constrained, possibly leading to an overestimation in stellar mass. Outshining and stochastic star formation are however still likely to have an effect on stellar masses derived from SED fitting, even when MIRI data is used (Narayanan et al., 2023; Shen et al., 2023), and the derived discrepancy in stellar mass without MIRI will likely depend on the assumed SFH, dust law and parameter priors, as we have shown they can also systematically change stellar mass estimates. Wang et al. (2024) also find that high mass galaxies at may require a higher star formation efficiency () than the local Universe, but they do not find any incompatibility with standard cosmological models.
@@
6.2 Comparison of our sample to other studies
The study by Labbé et al. (2023) found a collection of unexpectedly massive, high-redshift galaxies in the CEERS field. These objects typically have blue colors in the short wavelength NIRCam filters, but extremely red colors at longer wavelengths. We have cross-matched their 13 candidates to our catalogue in order to compare our derived masses and redshifts. In the following text we give both the ID in our catalogue followed by the ID given in Labbé et al. (2023) for the sources we discuss here. 11 of their candidates appear within our catalogue, of which 7 are selected by our selection criterion and are among the most massive galaxies we find. The two sources we do not recover appear to be blended with neighbouring sources in our SExtractor segmentation maps. Of the four that we detect but do not select, 5346_CEERSP6 (11184) falls within our image mask (close to the detector gap in the NIRCam SW observations), and the other three fail our signal to noise requirement in the rest-frame UV (5 sigma). One of these, 2683_CEERSP3 (13050), has been spectroscopically confirmed as a broadline AGN by Kocevski et al. (2023) at . We do not detect 37888 or 39575, which are among the lowest mass candidates presented in Labbé et al. (2023). We do not consider 2 of the selected galaxies in this study as their best-fitting redshift from EAZY-py is marginally below our z6.5 limit. For these galaxies, 8750_CEERSP3 (7274) and 2499_CEERSP1 (25666), we find lower redshift solutions than Labbé et al. (2023), with and vs their and . For four of the remaining five our photo- estimates agree well with Labbé et al. (2023), but for 1516_CEERSP2 (21834), we find a significantly higher photo- solution of with EAZY-py. However our fiducial Bagpipes results, which incorporate the EAZY-py photo- posterior as a prior, suggest a photo- closer to , which is in agreement with their result of . For these five galaxies, our stellar mass estimates are on average a factor of 2.1 (0.32 dex) smaller than Labbé et al. (2023), which is mostly due to the differing choice of IMF. Madau & Dickinson (2014) suggests a scaling factor of to convert a stellar mass estimate from Salpeter (1955) to Kroupa (2001), which does not account for the full discrepancy between our mass estimates.
The most massive candidate presented by Labbé et al. (2023), 7463_CEERSP1 (38094), with a stellar mass of is the second most massive galaxy in our sample, with a fiducial Bagpipes mass of . Given that our sample contains the area of the dataset used in Labbé et al. (2023), this may suggest that the CEERS field is overdense. As a comparison we compute the GSMF implied by the results of Labbé et al. (2023), assuming 100% completeness. The results of this are shown in Figure 8. Their implied datapoint at both redshift 8 and 9 lies above our best-fitting Schechter functions, due to a combination of a smaller cosmic volume and higher stellar masses than our results. Wang et al. (2024) shows that the stellar masses of high-z galaxies selected with the same color criteria as Labbé et al. (2023) are reduced by dex when rest-frame m observations from MIRI are included in SED fitting, suggesting that the stellar masses derived without MIRI data for the most massive galaxies in this work and in Labbé et al. (2023) may be systematically overestimated.
Although our source detection is done at the longest available wavelengths, our JWST sample is primarily selected in the rest-frame UV due to our Lyman-break criteria. Galaxies with weak UV emission may not be selected or even detected in our sample, whilst still containing significant stellar mass. Highly obscured galaxies like Submillimeter Galaxies (SMGs), which have high stellar masses, will not be included in our selection criteria, although they are thought to have a low spatial density (e.g. Smail et al., 2023).
Whilst high- galaxies are predicted to be relatively compact (Roper et al., 2022; Shen et al., 2024), the high resolution of JWST and relatively small extraction apertures used also means we are more sensitive to compact, high-surface brightness sources than extended, low-surface brightness sources. If a population of extended, diffuse galaxies exists at these redshifts they are likely to be excluded from this sample.
@@ 6.3 Contamination of the GSMF by hidden AGN
As discussed in § 5.2.3, obscured high- AGN can impact the high-mass end of the GSMF at by masquerading as galaxies with evolved stellar populations and therefore higher masses. Kokorev et al. (2024) presents candidate LRDs in both the CEERS and GOODS-South fields, which covers the NGDEEP and JADES-Deep observations. We compare their catalogue to our high- galaxy sample, finding 5(3) of their CEERS (GOODS-South) candidates in our robust sample. They are all at in our sample, with IDs 1080_CEERSP9, 2184_CEERSP1, 4842_CEERSP6, 7463_CEERSP1, 7520_CEERSP5, 6169_JADES-Deep-GS, 31738_JADES-Deep-GS and 36222_JADES-Deep-GS. It is worth noting that some of these galaxies in CEERS were first found in Labbé et al. (2023) and are discussed in § 6.2. We detect almost all of the galaxies from Kokorev et al. (2024) in the CEERS and GOODS-South fields (18 and 10 respectively), but for the galaxies that do not appear in our final sample we either find a lower redshift solution (below our cutoff), or we do not robustly detect the Lyman break to 5, meaning they do not meet our selection criteria. We see good agreement in redshift, with a maximum offset of , and a mean offset of . These galaxies are found to have high stellar masses in our SED-fitting, with a median stellar mass of = 10.30 in our fiducial Bagpipes results, with five of them forming the most massive galaxies in our fiducial sample. It is likely that at least some of these objects do contain AGN, which would lower the inferred stellar masses and hence lower the high-mass end of the GSMF. Further observations with NIRSpec or MIRI are essential to ascertain the true nature of these sources. In Appendix A we recompute the GSMF without any LRDs and show that these sources dominate the high mass end of the GSMF. D’Silva et al. (2023) has shown that accounting for the contribution of AGN lowers the cosmic star formation rate density by 0.4 dex at , which will also lower the inferred stellar mass density, and finds that a significant fraction of the LRDs are hidden AGN.

### v3

@@ 5.4 Fitting the GSMF
The Schechter function, which consists of a power-law with an exponential fall-off, is parameterised by 3 parameters: , M⋆, and . controls the slope of the low-mass end of the SMF, while M⋆ gives the turnover point at which the function turns from a power-law to an exponential fall-off. controls the overall scaling of the SMF, and is the only parameter that has been confidently shown to evolve with cosmic time (Popesso et al., 2023).
Before JWST the mass function was constrained up to using HST WFC3/IR, Spitzer/IRAC and ground-based data (Visible and Infrared Survey Telescope for Astronomy (VISTA), United Kingdom Infrared Telescope (UKIRT)) observations of the HUDF, CANDELS, Cluster Lensing and Supernova survey with Hubble (CLASH) and Hubble Frontier Fields (HFF) surveys (Stark et al., 2013; Bradač et al., 2014; Duncan et al., 2014; Bouwens et al., 2016; Song et al., 2016; Stefanon et al., 2017; Bhatawdekar et al., 2019; Kikuchihara et al., 2020). These studies generally agree that the mass function steepens out to , with the low-mass slope, , . Duncan et al. (2014); Bhatawdekar et al. (2019); Stefanon et al. (2021) find that the stellar mass to halo mass ratio in galaxies (M⋆-Mh) shows no evolution over despite a 3 dex increase in overall stellar mass, suggesting that stellar and halo mass grew together during reionization. Pre-JWST, the accuracy of stellar mass functions at these redshifts is limited by the uncertainty in stellar mass estimates from Spitzer data, where the broadband photometry can be affected by strong nebular emission causing scatter in stellar mass estimates, which is compounded by the limited number of spectroscopically confirmed galaxies at these redshifts (Roberts-Borsani et al., 2016; Strait et al., 2020). Previously, measurements of the mass function above have relied on UV-selected samples of galaxies, with mass indirectly inferred through a calibrated L-M⋆ relationship (Song et al., 2016; Harikane et al., 2016) rather than direct measurements of the rest-frame optical wavelengths. Photometric mass-selected samples above often have significant unavoidable mass uncertainties at higher redshift (Retzlaff et al., 2010; Laigle et al., 2016; Weaver et al., 2022), due to the lack of available rest optical or NIR photometry. Many studies combine ground and space-based observations in order to probe the entire mass function, although combining these datasets robustly is difficult due to systematics between different surveys, including differing selection functions, SED fitting techniques, detection bands, and survey depths.
@@
For our highest redshift bin (11.5 13.5), we do not attempt to fit the mass function for a number of reasons; firstly given the low number of available data point points a meaningful fit is not possible without fixing multiple parameters, and secondly due to the increasing uncertainties in stellar mass estimates and reliability of candidates at these redshifts. Furthermore the stellar mass estimates are less reliable when there is less rest-frame optical photometry available during SED fitting. Our JAGUAR simulations also suggest that this bin has high contamination at lower mass, with many galaxies below 108.5 M⊙ having 50% average contamination across the fields they are detectable in.
We use wide uniform priors on , M⋆ and , with and in our MCMC fitting procedure and run until it has converged before taking independent draws from the posterior. We calculate the median posterior, the uncertainty from the 16th/84th percentiles as well as the maximum likelihood draw.
@@
Figure 8 shows the covariance of the Schechter parameters for each GSMF fit. A range of parameters are able to fit our GSMF estimates, demonstrating the highly covariant nature of the Schechter parameters. We note that the confidence intervals for the turnover mass, M⋆, and the normalization, , appear somewhat unconstrained at the high and low ends respectively. The lack of constraint on M⋆ is a consequence of the wide stellar mass bins used and consequential low number of overall mass bins. Wherever we find M11.5, the GSMF is fit by a pure power-law only, and there is no reason to continually extend the prior outside the range where we have available data to constrain M⋆.
@@
Redshift Bin | Med() | Mpc-3) | Comp (%) | Cont (%) | Ngal | |
6.5 7.5 | 8.07 | 31 | 23 | 273 | ||
@@ 5.6 Stellar Mass Density
We perform the integration given in Equation 5, where is the best-fitting Schechter function fit for each redshift bin. We make use of the quad function from SciPy (Virtanen et al., 2020), integrating between 108 M⊙ and 1013 M⊙. We compare the integration of both the maximum likelihood fit as well as the median posterior fit. For comparison we also integrate the best-fitting Schechter functions from the literature in each redshift bin, converting to a Kroupa (2001) IMF if necessary.
@@ 6.1 Massive Galaxies in the Early Universe
Numerous studies have reported an excess of galaxies with high inferred stellar mass at high-redshift (Labbé et al., 2023; Endsley et al., 2023a; Akins et al., 2023; Xiao et al., 2023). If galaxies of the inferred masses do exist in the number densities implied by these studies, this could represent a challenge to CDM cosmology or our understanding of high-redshift galaxy astrophysics, given the available timescale and available gas reservoirs (Labbé et al., 2023; Lovell et al., 2023; Boylan-Kolchin, 2023).
@@
In order to test whether our derived stellar mass and redshift estimates are in tension with CDM we use the Extreme-Value Statistics (EVS) approach, presented in Lovell et al. (2023) and available online666github.com/christopherlovell/evstats. We follow the methodology of Lovell et al. (2023) throughout this section. EVS is a method for predicting the distribution of the most extreme values from a given distribution. In this case we use a parametrisation of the high-z halo mass function (Behroozi et al., 2013), and predict the PDF of the highest mass halo in some volume (given by a redshift interval, sky fraction and assumed cosmology).
Figure 11 shows these individual PDF contours projected into the mass-redshift plane. We show the 1, 2 and 3 sigma contours around the median prediction for the most massive galaxy at a given redshift with a dot-dash line and shaded contours. A unique aspect of the EVS approach is that it predicts both upper and lower bounds to the most massive object; the most massive galaxy can be too small for the assumed cosmology and astrophysics, as well as too massive. A universal baryon fraction of 0.16 is assumed based on cosmological results (Planck Collaboration et al., 2016), and we compute the EVS limits for a range of stellar fractions (fraction of baryons which form stars). In Figure 11 we show the 3 upper limit, assuming a stellar fraction of unity, with a solid black line.
@@
Overlaid on Figure 11 we show the SED-fitting derived redshifts and corrected stellar masses for our sample. We correct our derived stellar masses for Eddington bias (Eddington, 1913) following . is the stellar mass estimate, is the uncertainty in the stellar mass taken from the posterior PDF, and is the local slope of the halo mass function. Jespersen et al. (2024) looks at the impact of cosmic variance on the predictions of EVS, as this is not accounted for in the method of Lovell et al. (2023) we use here. Given that our sample consists of multiple widely-separated fields, we do not believe cosmic variance will play a significant role on this EVS analysis. No galaxies in our sample fall above the 3 limit given a stellar fraction of unity, shown with a black line, which would require more stellar mass than the available baryons to form stars. For galaxies with redshift and stellar mass estimates which place them in possible tension with the EVS limits assuming a realistic stellar fraction (more than 1.5 the mean value of the lognormal distribution, shown by the dashed-dot line in Figure 11) we show the individual mass measurements on the figure. Galaxies below this limit are not shown. For each of these galaxies we show the ‘fiducial’ Bagpipes photo- and stellar mass, as well as the maximum and minimum stellar mass estimate for that galaxy in our Bagpipes and Prospector fitting, including the modified IMF results we discuss in § 4.4. We link individual mass estimates for the same object with dotted lines. In the majority of cases, whilst the maximum stellar mass estimate may suggest a possible tension, the minimum and often the fiducial mass estimates are not in tension.
@@
It is worth noting that all galaxies which show a possible tension at redshift in this figure are classed as ‘little red dots’ (LRDs), which are discussed further in § 5.2.3 and potentially contain a significant AGN component we do not account for. The mass estimates for these galaxies may be overestimated by up to 1 dex (Williams et al., 2024), as it is impossible to disentangle the relative contributions of stellar and AGN emission to the broadband photometry, and difficult even with spectroscopy (e.g. Wang et al., 2024). It is clear from Figure 11 that none of our candidates are in tension with the predictions of CDM, as seen by the lack of sources above the solid black line, however a number of objects do require very high stellar fractions at these redshifts. The galaxies which require the highest star formation efficiencies seem to be found at , rather than the higher redshift probed in this study. This has also been found for HST-dark galaxies observed as part of FRESCO, where the galaxies with highest implied stellar mass densities are between (Xiao et al., 2023).
@@
Wang et al. (2024) have shown that stellar masses of high mass galaxies are typically overestimated by 0.4 dex when m rest-frame emission is not used in SED fitting, which requires MIRI observations in this redshift regime, and as much as 0.6 - 1 dex for the reddest sources (Williams et al., 2024).As MIRI observations are not used in this study, the degeneracies in age-attenuation and the relative contributions of strong emission lines and the stellar continuum observed by Wang et al. (2024) are not constrained, possibly leading to an overestimation in stellar mass. Outshining and stochastic star formation are however still likely to have an effect on stellar masses derived from SED fitting, even when MIRI data is used (Narayanan et al., 2023; Shen et al., 2023), and the derived discrepancy in stellar mass without MIRI will likely depend on the assumed SFH, dust law and parameter priors, as we have shown they can also systematically change stellar mass estimates. Wang et al. (2024) also find that high mass galaxies at may require a higher star formation efficiency () than the local Universe, but they do not find any incompatibility with standard cosmological models. Alternative explanations which do not require high SFE, such as a blue-tilted primordial power spectrum, have also been proposed in the literature (e.g. Parashari & Laha, 2023).
@@ 6.2 Contamination of the GSMF by hidden AGN
The study by Labbé et al. (2023) discovered unexpectedly massive, high-redshift galaxies in the CEERS field known as ’Little Red Dots’, as discussed in § 5.2.3. Kokorev et al. (2024) used similar selection criteria to extend this to the GOODS-S field. In this section we compare our photometric redshifts and mass estimates and investigate the impact on the GSMF. We cross-matched their 13 candidates with our catalogue to compare masses and redshifts. Of these, 11 are in our catalogue, 7 of which meet our selection criteria and are among the most massive galaxies we identified. Two candidates were not recovered due to blending with neighboring sources in our SExtractor segmentation maps. Four others we detected but did not select: 5346_CEERSP6 (11184) is near a detector gap, and the other three did not meet our UV signal-to-noise requirement. Notably, 2683_CEERSP3 (13050) is a spectroscopically confirmed broadline AGN at (Kocevski et al., 2023).
We did not detect 37888 or 39575, the lowest mass candidates in Labbé et al. (2023) as they were blended with neighbours. Two galaxies were excluded from our study as their redshifts were below our limit. For these, we found lower redshifts than Labbé et al. (2023): 8750_CEERSP3 (7274) at and 2499_CEERSP1 (25666) at . For five galaxies, our photo- estimates matched well, except for 1516_CEERSP2 (21834), where our estimate was with EAZY-py, differing to Bagpipes which found , aligning with their result of . Our stellar mass estimates averaged 2.1 smaller than Labbé et al. (2023), largely due to differing IMFs. The most massive candidate, 7463_CEERSP1 (38094), with a stellar mass of , is the second most massive in our sample, with a Bagpipes mass of . Given our larger survey area, this suggests the CEERS field might be overdense, as suggested by Desprez et al. (2024). We computed the GSMF implied by Labbé et al. (2023), showing their datapoints at and are above our best-fitting Schechter functions, due to smaller cosmic volume and higher stellar masses. Wang et al. (2024) shows that including MIRI data in SED fitting reduces high-z galaxy masses by dex, indicating potential overestimation without MIRI data. When comparing to Labbé et al. (2023), we find 8 of their CEERS & GOODS-South candidates in our robust sample, with 28 more being detected but below our cutoff or below our rest UV SNR requirements. We see good agreement in redshift, with a maximum offset of , and a mean offset of . These galaxies are found to have high stellar masses in our SED-fitting, with a median stellar mass of = 10.30 in our fiducial Bagpipes results, with five of them forming the most massive galaxies in our fiducial sample. They are all at in our sample. Further observations with NIRSpec or MIRI are essential to ascertain the true nature of these sources. In Appendix B we recompute the GSMF without any LRDs and show that these sources dominate the high mass end of the GSMF. D’Silva et al. (2023) has shown that accounting for the contribution of AGN lowers the cosmic star formation rate density by 0.4 dex at , which will also lower the inferred stellar mass density, and finds that a significant fraction of the LRDs are hidden AGN.
Our sample is primarily selected in the rest-frame UV, potentially missing galaxies with weak UV emission but significant stellar mass, such as Submillimeter Galaxies (SMGs). These high-mass but low spatial density galaxies are excluded from our selection. Additionally, our high-resolution detection favors compact, high-surface brightness sources over extended, low-surface brightness ones, possibly excluding a population of extended, diffuse galaxies at these redshifts.